1. All of the following groups can be hurt by unanticipated inflation, except:

a- Lenders

b- Borrowers

c- Retired People on Fixed Incomes

d- Savers

2. The type of inflation that occurs because input prices (costs of production) have risen rapidly is:

a- Deflation

b- Demand-Pull Inflation

c- Cost-Push Inflation

d- Hyperinflation

3. Suppose that the CPI in 2000 was 176 and the CPI in 2001 was 178.4. Calculate the rate of inflation from 2000 to 2001.

a- 2.4%

b- 5.6%

c- 1.36%

d- 1.04%

4. Suppose that the value of the CPI basket in 2010 was $36,500 and the value of the same basket in the base year (1982-84) was $20,000. What is the value of the CPI in 2010?

a- 1.825

b- 182.5

c- 5.47

d- 54.7

5. Suppose that an economy has 9 million people working full-time. It also has 1 million people who are actively seeking work but currently unemployed as well as 2 million discouraged workers who have given up looking for work and are currently unemployed. What is this economy’s unemployment rate?

a- 15%

b- 25%

c- 20%

d- 10%

6. Suppose that in a population of 400,000 people, there are 100,000 people who are either institutionalized or are too young to work. Additionally, there are 60,000 people who choose not to work because they are stay-at-home parents, college students, or retired. In this scenario, if 228,000 people currently are employed full-time, what is the unemployment rate?

a- 2.28%

b- 57%

c- 10%

d- 5%

7. Suppose that in a population of 400,000 people, there are 100,000 people who are either institutionalized or are too young to work. Additionally, there are 60,000 people who choose not to work because they are stay-at-home parents, college students, or retired. In this scenario, what is the size of the labor force?

a- 300,000

b- 400,000

c- 240,000

d- 360,000

8. All of the following are causes of shocks that can cause business cycles, except:

a- Political Events

b- Sudden or Dramatic Changes in Productivity

c- Consistent, Steady Innvation

d- Financial Instability

9. Which is the correct sequencing of the business cycle?

a- Expansion, Recession, Trough, Peak

b- Peak, Expansion, Recession, Trough

c- Expansion, Peak, Recession, Trough

d- Trough, Recession, Expansion, Peak

10. Economists define a recession as:

a- Two or more consecutive quarters or declining real GDP

b- Any reduction in employment that occurs from one period to the next

c- Whenever GDP from one period is lower than the period before it

d- Six months of declining GDP
Answer #1

1. Option B. Borrowers

Explanation: Borrowers are benefited by unanticipated inflation as they need to return money worth less than what was borrowed.

2. Option C

Explanation: Cost-push inflation takes place with an increase in the production cost.

3. Option C

Explanation: Rate of inflation = % change in CPI from 2000 to 2001 = (178.4 - 176)/176 * 100 = 2.4/176 * 100 = 240/176 = 1.36%

4. Option B

Explanation: The % change in the value of CPI basket = (36,500 - 20,000)/20,000 * 100 = 16,500/20,000 * 100 = 165/200 *100 = 0.825 * 100 = 82.5%

So, CPI would rise 82.5% from 100. So, CPI in 2010 = 182.5